### v1

# Robust Weight Initialization for Tanh Neural Networks with Fixed Point Analysis

Abstract
As a neural network’s depth increases, it can achieve strong generalization performance. Training, however, becomes challenging due to gradient issues. Theoretical research and various methods have been introduced to address this issues. However, research on weight initialization methods that can be effectively applied to tanh neural networks of varying sizes still needs to be completed. This paper presents a novel weight initialization method for Feedforward Neural Networks with tanh activation function. Based on an analysis of the fixed points of the function $\tanh(ax)$, our proposed method aims to determine values of $a$ that prevent the saturation of activations. A series of experiments on various classification datasets demonstrate that the proposed method is more robust to network size variations than the existing method. Furthermore, when applied to Physics-Informed Neural Networks, the method exhibits faster convergence and robustness to variations of the network size compared to Xavier initialization in problems of Partial Differential Equations.
1 Introduction
Deep learning has enabled substantial advancements in state-of-the-art performance across various domains (LeCun et al., 2015; He et al., 2016). In general, the expressivity of neural networks exponentially increases with depth (Poole et al., 2016; Raghu et al., 2017), enabling strong generalization performance. This increased depth, though, can result in vanishing or exploding gradients and poor signal propagation throughout the model (Bengio et al., 1993), prompting the development of various weight initialization methods. Xavier initialization (Glorot & Bengio, 2010) ensures signals stay in the non-saturated region for sigmoid and hyperbolic tangent activations, while He initialization (He et al., 2015) maintains stable variance for ReLU networks. Especially in ReLU neural networks, several weight initialization methods have been proposed to mitigate the dying ReLU problem, which hinders signal propagation in deep networks (Lu et al., 2019; Lee et al., 2024). However, to the best of our knowledge, research on the initialization method to tackle the stability of extremely deep tanh networks during training is still limited. Such networks commonly use Xavier initialization (Raissi et al., 2019; Jagtap et al., 2022; Rathore et al., 2024) and are widely applied in various domains, such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and Recurrent Neural Networks (RNNs) (Rumelhart et al., 1986), with performance often dependent on model size and initialization randomness (Liu et al., 2022).
The main contribution of this paper is the proposal of a simple weight initialization method for FeedForward Neural Networks (FFNNs) with tanh activation function. This method facilitates effective learning across a range of network sizes, outperforming Xavier initialization by reducing the need for extensive hyperparameter tuning such as the number of hidden layers and units. The theoretical foundation for this approach is provided through the fixed point of the function $\tanh(ax)$. We experimentally demonstrate that the proposed method achieves higher validation accuracy and lower validation loss compared to the Xavier initialization method across various FFNN network sizes on the MNIST, Fashion MNIST, and CIFAR-10 datasets. Additionally, the proposed method demonstrates its effectiveness in training across various network configurations within PINNs. Notably, while Xavier initialization shows decreasing loss as network depth increases, it fails to maintain performance beyond a certain depth, leading to increased loss and poor training outcomes. In contrast, the proposed method continues to improve performance even at greater depths, ensuring stable training and better results.
Contributions. Our contributions can be summarised as follows:
  • We show the conditions under which activation values do not vanish as we increase the depth of the neural network, using a fixed-point analysis (Section 3.1 and 3.2).
  • We propose a novel weight initialization method for tanh-based neural networks that has strong robustness to variations in network size (Section 3.2 and 3.3).
  • We experimentally demonstrate that the proposed method is more robust to variations in network size than Xavier initialization on image benchmark datasets and PINNs (Section 4).
2 Related works
The expressivity of neural networks typically grows exponentially with depth, resulting in improved generalization performance (Poole et al., 2016; Raghu et al., 2017). Weight initialization is crucial for training deep networks effectively (Saxe et al., 2014; Mishkin & Matas, 2016). Xavier (Glorot & Bengio, 2010) and He He et al. (2015) initialization are common initialization methods typically used with tanh and ReLU activation functions, respectively. Various initialization methods have been proposed to facilitate the training of deeper ReLU neural networks (Lu et al., 2019; Bachlechner et al., 2021; Zhao et al., 2022; Lee et al., 2024). However, to the best of our knowledge, research on weight initialization for neural networks with tanh activation remains limited. Tanh neural networks have been increasingly used, particularly in physics-informed neural networks (PINNs).
PINNs have shown promising results in solving forward, inverse, and multiphysics problems arising in science and engineering. (Lu et al., 2021; Karniadakis et al., 2021; Cuomo et al., 2022b; a; Yin et al., 2021; Wu et al., 2023; Hanna et al., 2022; Bararnia & Esmaeilpour, 2022; Shukla et al., 2020; Zhu et al., 2024; Hosseini et al., 2023; Mao et al., 2020). PINNs approximate solutions to partial differential equations (PDEs) using neural networks and are typically trained by minimizing a loss defined by the sum of least-squares that incorporates the residual of PDE, boundary conditions, and initial conditions. This loss is usually minimized using gradient-based optimizers such as Adam (Kingma, 2014), L-BFGS (Liu & Nocedal, 1989), or a combination of both. Universal approximation theories (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Park et al., 2020; Guliyev & Ismailov, 2018b; Shen et al., 2022; Guliyev & Ismailov, 2018a; Maiorov & Pinkus, 1999; Yarotsky, 2017; Gripenberg, 2003) guarantee the capability and performance of neural networks as an approximation of the analytic solution to PDE. However, PINNs still face challenges in accuracy, stability, computational complexity, and tuning optimal hyperparameters of loss terms. To alleviate these issues, many authors have introduced enhanced versions of PINNs: (1) the self-adaptive loss balanced PINNs (lbPINNs) that automatically adjust the hyperparameters of loss terms during the training process (Xiang et al., 2022), (2) the Bayesian PINNs (B-PINNs) that are specialized to deal with forward and inverse nonlinear problems with noisy data (Yang et al., 2021), (3) Rectified PINNs (RPINNs) that are trained with the gradient information from the numerical solution by the multigrid method and designed for solving stationary PDEs  (Peng et al., 2022), (4) Auxiliary Pinns (A-PINNs) that effectively handle integro-differential equations (Yuan et al., 2022), (5) conservative PINNs (cPINNs) and exetended PINNs (XPINNs) that adopt the domain decomposition technique (Jagtap et al., 2020; Jagtap & Karniadakis, 2020), (6) parrel PINNs that reduces the computational cost of cPINNs and XPINNs (Shukla et al., 2021), (7) gradient-enhanced PINNs (gPINNs) that use the gradient of the PDE loss term with respect to the network inputs  (Yu et al., 2022).
PINNs primarily employ Xavier initialization for training (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023; Song et al., 2024), but our experimental results indicate that this method limits the performance of larger network sizes. Although there have been recent results on initialization methods for PINNs, most of them have relied on transfer learning (Tarbiyati & Nemati Saray, 2023). Thus, we propose a weight initialization method that does not require transfer learning and is robust to variations in network size.
(a) Xavier Initialization
(b) Proposed Initialization
Figure 1: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using Xavier initialization (Left) and the proposed initialization (Right). Experiments were conducted on networks with 10,000 hidden layers, each having the same number of nodes: $16$, $32$, $64$, or $128$.
3 Proposed Weight Initialization method
In this section, we discuss the proposed weight initialization method. Subsection 3.1 introduces the theoretical motivation behind the methodology. Subsection 3.2 presents how to derive the initial weight matrix that satisfies the conditions outlined in Subsection 3.1. Finally, in Subsection 3.3, we suggest the optimal hyperparameter $\sigma_{z}$ in the proposed method.
3.1 Theoretical motivation
Experimental results in Figure 1 reveal that when Xavier initialization is employed in FFNNs with tanh activation, the distribution of activation values tends to cluster around zero in deeper layers. This vanishing of activation values can hinder the training process due to a discrepancy between the activation values and the desired output. However, theoretically preventing this phenomena is not straightforward. In this subsection, we gives a theoretical analysis based on a fixed point of $\tanh(ax)$ to bypass the phenomena. Before giving the theoretical foundations, consider the basic results for a tanh activation function. Recall that $x^{*}$ is a fixed point of a function $f$ if $x^{*}$ belongs to both the domain and the codomain of $f$, and $f(x^{*})=x^{*}$. The proofs of Lemma 1 and Lemma 2 are provided in Appendix A.
Lemma 1.
For a fixed $a>0$ define the function $\phi_{a}:\mathbb{R}\to\mathbb{R}$ given as
$\phi_{a}(x):=\tanh(ax).$
Then, there exists a fixed point $x^{\ast}$. Furthermore,
  • $(1)$
    if $0<a\leq 1$, then $\phi$ has a unique fixed point $x^{\ast}=0$.
  • $(2)$
    if $a>1$, then $\phi$ has three distinct fixed points: $x^{\ast}=-\xi_{a}$, 00, $\xi_{a}$ such that $\xi_{a}>0$.
Remark that the function $\phi_{a}$ can be considered as one-layer tanh FFNN.
Lemma 2.
For a given initial value $x_{0}>0$ define
$x_{n+1}=\phi_{a}(x_{n}),\quad n=0,1,2,\ldots.$
Then $\{x_{n}\}_{n=1}^{\infty}$ converges regardless of the positive initial value $x_{0}>0$. Moreover,
  • $(1)$
    if $0<a\leq 1$, then $x_{n}\to 0$ as $n\to\infty$.
  • $(2)$
    if $a>1$, then $x_{n}\to\xi_{a}$ as $n\to\infty$.
Note that the parameter $a$ in Lemma 2 does not change across all iterations. In Propositions 3 and Corollary 4, we address cases where the value of $a$ varies with each iteration.
Proposition 3.
Let $\{a_{n}\}_{n=1}^{\infty}$ be a positive real sequence, i.e., $a_{n}>0$ for all $n\in\mathbb{N}$, such that only finitely many elements are greater than $1$. Suppose that $\{\Phi_{m}\}_{m=1}^{\infty}$ is a sequence of functions defined as for each $m\in\mathbb{N}$
$\Phi_{m}=\phi_{a_{m}}\circ\phi_{a_{m-1}}\circ\cdots\circ\phi_{a_{1}}.$
Then for any $x\in\mathbb{R}$
$\lim_{m\to\infty}\Phi_{m}(x)=0.$
Proof.
Set $N=\max\{n|a_{n}>1\}$. Define the sequences $\{b_{n}\}_{n=1}^{\infty}$ and $\{c_{n}\}_{n=1}^{\infty}$ such that $b_{n}=c_{n}=a_{n}$ for $n\leq N$, with $b_{n}=0$ and $c_{n}=1$ for $n>N$. Suppose that $\{\hat{\Phi}_{m}\}_{m=1}^{\infty}$ and $\{\tilde{\Phi}_{m}\}_{m=1}^{\infty}$ are sequences of functions defined as for each $m\in\mathbb{N}$
$\hat{\Phi}_{m}=\phi_{b_{m}}\circ\phi_{b_{m-1}}\circ\cdots\circ\phi_{b_{1}},\quad\tilde{\Phi}_{m}=\phi_{c_{m}}\circ\phi_{c_{m-1}}\circ\cdots\circ\phi_{c_{1}}.$
Then, the inequality $\hat{\Phi}_{m}\leq\Phi_{m}\leq\tilde{\Phi}_{m}$ holds for all $m$. By Lemma 1, for any $x\geq 0$, we have $\lim_{m\to\infty}\hat{\Phi}_{m}=0$ and $\lim_{m\to\infty}\tilde{\Phi}_{m}=0$. Therefore, the Squeeze Theorem guarantees that $\lim_{m\to\infty}\Phi_{m}(x)=0$.
Corollary 4.
Let $\epsilon>0$ be given. Suppose that $\{a_{n}\}_{n=1}^{\infty}$ be a positive real sequence such that only finitely many elements are lower than $1+\epsilon$. Then for any $x\in\mathbb{R}\setminus\{0\}$
$\left|\lim_{m\to\infty}\Phi_{m}(x)\right|\geq\xi_{1+\epsilon}$
Proof.
Set $N=\max\{n\mid a_{n}<1+\epsilon\}$. Define the sequence $\{b_{n}\}_{n=1}^{\infty}$ such that $b_{n}=a_{n}$ for $n\leq N$, and $b_{n}=1+\epsilon$ for $n>N$. The remainder of the proof is analogous to the proof of Proposition 3. ∎
3.2 The derivation of the proposed weight initialization method
To establish the notation, consider a feedforward neural network with $L$ layers. The network processes $K$ training samples, denoted as pairs $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{K}$, where $\bm{x}_{i}\in\mathbb{R}^{N_{x}}$ is training input and $\bm{y}_{i}\in\mathbb{R}^{N_{y}}$ is its corresponding output. The iterative computation at each layer $\ell$ is defined as follows:
$\bm{x}^{\ell}=\tanh(\mathbf{W}^{\ell}\bm{x}^{{\ell}-1}+\mathbf{b}^{\ell})\in\mathbb{R}^{N_{\ell}}\quad\text{for all }\ell=1,\ldots,L,$
where $\mathbf{W}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ is the weight matrix, $\mathbf{b}^{\ell}\in\mathbb{R}^{N_{\ell}}$ is the bias, and $\tanh(\cdot)$ is an element-wise activation hyperbolic tangent function.
We present a simplified analysis of signal propagation in FFNNs with the tanh activation function. For notational convenience, it is assumed that all hidden layers, as well as the input and output layers, have a dimension of $n$, i.e., $N_{\ell}=n$ for all $\ell$. Given an arbitrary input vector $\bm{x}=(x_{1},\ldots,x_{n})$, the first layer activation $\bm{x}^{1}=\tanh(\mathbf{W}^{1}\bm{x})$ can be expressed component-wise as:
$x^{1}_{i}=\tanh\big{(}w^{1}_{i1}x_{1}+\cdots+w^{1}_{in}x_{n}\big{)}=\tanh\bigg{(}\Big{(}w^{1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{1}_{ij}x_{j}}{x_{i}}\Big{)}x_{i}\bigg{)},\,\text{for }i=1,\dots,n.$
For the $k+1$-th layer, $i=1,\ldots,n$, this expression can be generalized as:
$x_{i}^{k+1}=\tanh\left(a_{i}^{k+1}x_{i}^{k}\right),\,\text{where}\quad a_{i}^{k+1}=w^{k+1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{k+1}_{ij}x_{j}^{k}}{x_{i}^{k}}.$ (1)
According to Lemma 2, when $a>1$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$, the sequence $\{x_{k}\}$ defined by $x_{k+1}=\tanh(ax_{k})$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $k\to\infty$. This result indicates that the sequence converges to the fixed point $\xi_{a}$ regardless of the initial value $x_{0}$ and ensures that the activation values do not vanish as network depth increases. Furthermore, by Lemma 2, if $a^{k}_{i}\leq 1$ for all $N\leq k\leq L$, then $x_{i}^{L}$ approaches zero. Therefore, to ensure that ((((i)))) $a_{i}^{k}$ remains close to 1 and ((((ii)))) $a_{i}^{k}\leq 1$ does not hold for all $N\leq k\leq L$, we design the initial weight matrix as $\mathbf{W}^{\ell}=\mathbf{D}^{\ell}+\mathbf{Z}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$, where $\mathbf{D}^{\ell}_{i,j}=1$ if $i\equiv j\pmod{N_{\ell-1}}$, 0 otherwise, and $\mathbf{Z}^{\ell}$ is a noise matrix drawn from $\mathcal{N}(0,\sigma_{z}^{2})$, where $\sigma_{z}$ is set to $\alpha/\sqrt{N^{\ell-1}}$ and $\alpha=0.085$. Then $a_{i}^{k+1}$ follows the distribution:
$a_{i}^{k+1}\sim\mathcal{N}\Bigg{(}1,\,\sigma_{z}^{2}+\sigma_{z}^{2}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\bigg{(}\frac{x_{j}^{k}}{x_{i}^{k}}\bigg{)}^{2}\Bigg{)}.$ (2)
According to Equation 2, the mean of $a_{i}^{k+1}$ is $1$, so choosing an appropriate $\sigma_{z}$ satisfies condition ((((i)))). For condition ((((ii)))), if $x_{i}^{k}$ becomes small relative to other elements in $\bm{x}^{k}$, the variance of $a_{i}^{k+1}$ increases, as indicated by Equation 2. As a result, the probability that the absolute value of $x_{i}^{k+1}$ surpasses that of $x_{i}^{k}$ is higher. However, if $\sigma_{z}$ is too small, the increase in the variance of $a_{i}^{k+1}$ becomes limited. Therefore, choosing an appropriate $\sigma_{z}$ is crucial.
(a) Same dimension
(b) Varying dimensions
Figure 2: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using the proposed initialization with $\alpha$ set to $0.04,0.085,0.15,$ and $0.5$. Network with $10,000$ hidden layers, each with 32 nodes (Left), and a network with alternating hidden layers of $64$ and $32$ nodes (Right).
3.3 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
In this subsection, we discuss how $\sigma_{z}$ impacts the scale of the activation values. Equation 2 indicates that $a_{i}^{k}$ follows a normal distribution, with variance depending on $\sigma_{z}$. Firstly, we experimentally investigated the impact of $\sigma_{z}$ on the scale of the activation values. As demonstrated in Figure 2, increasing $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ causes the activation values in any layer to be distributed over a broader range. However, setting $\sigma_{z}$ to a large value can lead to saturation, where most activations converge towards $-1$ and $1$. If $\sigma_{z}$ is too large, the probability that $a_{i}^{(k)}$ takes values far from $1$ (e.g., -10, 5, etc.) increases. This, in turn, increases the value of $1+\epsilon$ mentioned in Corollary 4, potentially bounding the activation values in sufficiently deep layers by $\xi_{1+\epsilon}$. Consequently, the activation values in deeper layers become less likely to approach zero and tend to saturate toward specific values. On the other hand, if $\sigma_{z}$ is too small, as mentioned in Subsection 3.2, the variance of $a_{i}^{k}$ becomes restricted. This is demonstrated experimentally in Figure 2, when $\alpha=0.00001$. For this reason, we experimentally found an optimal $\sigma_{z}=\alpha/\sqrt{N^{\ell-1}}$, with $\alpha=0.085$, that is neither too large nor too small. Results from experiments solving the Burgers’ equation using PINNs with varying $\sigma_{z}$ are presented in Appendix B.1.
Table 1: Validation accuracy and loss are presented for FFNNs with varying numbers of nodes $(2,8,32,128)$, each with $20$ hidden layers using the tanh activation function. All models were trained for $20$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented. When comparing different initialization methods under the same experimental settings, the better-performing method is highlighted in bold. Underlined values indicate the highest accuracy when only the number of nodes is varied.
2 8 32 128
MNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 49.78 1.632 68 0.958 91.67 0.277 95.45 0.154
Proposed 62.82 1.185 77.95 0.706 92.51 0.255 96.12 0.134
FMNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 42.89 1.559 68.55 0.890 81.03 0.533 86.2 0.389
Proposed 51.65 1.324 71.31 0.777 83.06 0.475 87.12 0.359
CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 32.82 1.921 43.51 1.608 48.62 1.473 47.58 1.510
Proposed 38.16 1.780 47.04 1.505 48.80 1.463 48.51 1.471
Table 2: Validation accuracy and loss are presented for FFNNs with varying numbers of layers $(10,50,100)$, each with $64$ number of nodes using the tanh activation function. All models were trained for $40$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented.
10 50 100
MNIST Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 96.55 0.112 96.57 0.123 94.08 0.194
Proposed 97.04 0.102 96.72 0.109 96.06 0.132
FMNIST Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 88.73 0.319 87.72 0.344 83.41 0.463
Proposed 89.42 0.305 88.51 0.324 86.01 0.382
CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 48.39 1.468 47.87 1.474 46.71 1.503
Proposed 48.41 1.458 48.71 1.461 48.96 1.437
4 Experiments
In this section, we conduct a series of experiments to validate the proposed weight initialization method. In Subsection 4.1, we evaluate the performance of an FFNN with the tanh activation function on benchmark datasets. In Subsection 4.2, we solve the Burgers’ equation and Allen-Cahn equation using Physics-Informed Neural Networks. Both experiments are conducted across various network sizes to verify whether the proposed method consistently performs well, independent of network depth and width. The experiments were conducted in TensorFlow without skip connections, normalization layers, and learning rate decay in any of the experiments.
4.1 Width Independence in Classification Task
To evaluate the effectiveness of the proposed weight initialization method, we conduct experiments on the MNIST, Fashion MNIST, and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, utilizing the Adam optimizer. All experiments are conducted with a batch size of 64 and a learning rate of 0.0001. Fifteen percent of the total dataset is allocated for validation.
We apply the proposed weight initialization method to evaluate its effectiveness in training tanh FFNNs, emphasizing its robustness to variations in network width. Four tanh FFNNs are created, each with 20 hidden layers, and with 2, 8, 32, and 128 nodes per hidden layer, respectively. In Table 1, for both the MNIST and Fashion MNIST datasets, the network with 128 nodes achieves the highest accuracy and lowest loss when our proposed method is employed. However, for the CIFAR-10 dataset, the network with 32 nodes yields the highest accuracy and lowest loss when employing the proposed method. In summary, our proposed method demonstrates robustness regardless of the number of nodes in tanh FFNNs. We provide more detailed experimental results in Appendix B.2.
4.2 Depth Independence in Classification Task
It is well known that the expressivity of neural networks generally increases exponentially with depth, enabling strong generalization performance (Poole et al., 2016; Raghu et al., 2017). Therefore, we employ the proposed weight initialization method to investigate its effectiveness in training deep FFNNs with the tanh activation function, emphasizing its robustness to variations in network depth. We create three tanh FFNNs, each with 64 nodes in all hidden layers, but with 10, 50, and 100 hidden layers, respectively. In Table 2, for both the MNIST and Fashion MNIST datasets, the network with 10 hidden layers achieves the highest accuracy and lowest loss when our proposed method is employed. Both initialization methods perform best in networks with the fewest layers, with performance degrading as the depth increases. However, for the CIFAR-10 dataset, we observe that the performance of the proposed method improves as the number of layers increases.
Furthermore, we conduct experiments with varying hidden layer dimensions, as shown in Figure 3. The network consists of 60 hidden layers, where the number of nodes alternates between 32 and 16 in each layer. We demonstrate superior performance in terms of both loss and accuracy across all epochs on the MNIST and CIFAR-10 datasets.
(a) MNIST Accuracy
(b) MNIST Loss
(c) CIFAR10 Accuracy
(d) CIFAR10 Loss
Figure 3: Validation accuracy and loss for a tanh FFNN with 60 hidden layers, where the number of nodes alternates between 32 and 16 across layers, repeated 30 times. The model was trained for 20 epochs on the MNIST and CIFAR-10 datasets.
4.3 Network size independence in pinn
Xavier initialization is the primary method used for training PINNs  (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023). In this section, we experimentally demonstrate that the method’s training performance is highly dependent on randomness and network size. Additionally, empirical results are provided demonstrating that the proposed method is more robust to variations in network size.
All experiments on Physics-Informed Neural Networks (PINNs) use full-batch training with a learning rate of 0.001. In this section, we solve the Allen-Cahn and Burgers’ equations using a tanh FFNN-based PINN with 20,000 collocation points. For the Allen-Cahn equation, the diffusion coefficient is set to $d=0.01$. The initial condition is defined as $u(x,0)=x^{2}\cos(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=-1$ and $u(1,t)=-1$, applied over the time interval $t\in[0,1]$. Similarly, for the Burgers’ equation, a viscosity coefficient of $\nu=0.01$ is employed. The initial condition is given by $u(x,0)=-\sin(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=0$ and $u(1,t)=0$ imposed for $t\in[0,1]$.
The Allen-Cahn equation is expressed as:
$\frac{\partial u}{\partial t}-d\frac{\partial^{2}u}{\partial x^{2}}=-\frac{u^{3}+u}{d}$
where $u(x,t)$ represents the solution, $d$ is the diffusion coefficient, and the nonlinear term $u^{3}-u$ models the phase separation dynamics.
The Burgers’ equation is given by:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}}$
where $u(x,t)$ is the velocity field, and $\nu$ is the viscosity coefficient.
Eight tanh FFNNs are created, each with 16 nodes in all hidden layers, but with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively. As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 20. However, as the depth increases, the loss gradually rises. In contrast, the proposed method achieves the lowest loss at a depth of 50 and maintains a loss of 0.00057 even at a depth of 80 layers. For the Burgers’ equation, the proposed method achieves the lowest loss at a depth of 60, while at the same depth, Xavier initialization results in a loss that is an order of magnitude higher (approximately $10^{2}$ difference).
Next, we double the number of nodes to observe the impact of node size on the loss. Eight new tanh FFNNs are created, each with 32 nodes in all hidden layers, and with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively. As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 30. However, as the depth increases, the model becomes untrainable, with a loss of 0.694 at a depth of 80. In contrast, the proposed method achieves the lowest loss at a depth of 40 and maintains a loss of 0.00059 even at a depth of 80 layers. For the Burgers’ equation, both methods show similar loss values up to a depth of 30. Beyond a depth of 40, however, the loss steadily increases with the Xavier method, while the proposed method records the lowest loss at a depth of 50.
Table 3: A PINN loss is presented for FFNNs with varying numbers of layers $(5,10,20,30,40,50,60,80)$ using the tanh activation function. The top table shows results with $16$ nodes per layer, and the bottom table shows results with $32$ nodes per layer. All models were trained for $300$ iterations using Adam and $300$ iterations using L-BFGS. The median PINN loss from the final iteration for the Burgers and Allen–Cahn equations, computed over 5 runs, is presented.
Allen-Cahn (16 Nodes) 5 10 20 30 40 50 60 80
Xavier 9.58e-04 8.16e-04 7.61e-04 1.06e-03 1.1e-03 1.24e-03 3.55e-03 1.81e-03
Proposed 9.21e-04 7.29e-04 5.76e-04 5.29e-04 5.37e-04 4.03e-04 4.73e-04 5.77e-04
Burgers (16 Nodes) 5 10 20 30 40 50 60 80
Xavier 6.97e-03 1.11e-02 7.9e-03 9.71e-03 2.45e-02 2.65e-02 6.5e-02 5.71e-02
Proposed 6.19e-03 5.08e-03 5.28e-03 9.31e-04 3.56e-03 8.27e-04 3.43e-04 2.05e-03
Allen-Cahn (32 Nodes) 5 10 20 30 40 50 60 80
Xavier 3.13e-01 5.03e-02 3.64e-03 2.37e-03 4.03e-03 5.27e-03 1.73e-02 6.94e-01
Proposed 1.04e-03 6.92e-04 5.34e-04 4.26e-04 3.31e-04 3.52e-04 3.85e-04 5.96e-04
Burgers (32 Nodes) 5 10 20 30 40 50 60 80
Xavier 1.12e-02 3.53e-03 2.72e-03 1.81e-03 7.60e-03 8.56e-03 9.86e-03 1.66e-01
Proposed 4.14e-03 4.11e-03 1.58e-03 1.29e-03 7.96e-04 5.85e-04 9.80e-04 1.47e-03
5 Conclusion
In this paper, we have introduced a novel weight initialization method for tanh FFNNs, grounded in the theoretical analysis of fixed points of the $\tanh(ax)$ function. Through our fixed-point analysis, we established conditions under which the vanishing or exploding of activation values can be prevented, even as the depth of the network increases.
Our proposed method exhibits strong robustness to variations in network size, as demonstrated across a variety of FFNN configurations and benchmark datasets, including MNIST, Fashion MNIST, and CIFAR-10. In contrast to Xavier initialization, which struggles to maintain stable performance as network depth increases, the proposed method consistently achieves superior results by preserving activation values. Furthermore, we explored the impact of the initialization hyperparameter $\sigma_{z}$ on the distribution of activation values. We demonstrated both theoretically and experimentally that the choice of $\sigma_{z}$ plays a significant role in maintaining the proper range of activations, balancing between vanishing and saturation. In the context of PINNs, the proposed initialization method shows improved performance in solving PDEs such as the Burgers’ equation and the Allen-Cahn equation. By maintaining a stable loss function and achieving faster convergence compared to Xavier initialization, our method demonstrates its practical utility in training networks for physical systems.
A key advantage of the proposed method lies in its robustness to network depth and width, significantly reducing the need for extensive hyperparameter tuning. By maintaining stable performance across varying network configurations, our approach helps to minimize the time and effort spent on searching for optimal network architectures, allowing researchers to focus on model design and other aspects of the training process. This makes the proposed method particularly valuable in large-scale and resource-constrained applications where efficient training is critical.
Appendix A Proofs of the theoretical results
A.1 Proof of Lemma 1
Proof.
We define $g(x)=\tanh(ax)-x$. Since $g(x)$ is continuous, and $g(-M)>0$, $g(M)<0$ for a large real number $M\in\mathbb{R}^{+}$, the Intermediate Value Theorem guarantees the existence of a point $x$ such that $g(x)=0$.
First, consider the case $0<a\leq 1$. Since $0<a\leq 1$, the derivative $g^{\prime}(x)=a\cdot\text{sech}^{2}(ax)-1$ satisfies $-1\leq g^{\prime}(x)\leq a-1<0$ for all $x$. Hence, $g(x)$ is strictly decreasing and therefore $g(x)$ has the unique root. At $x=0$, $\phi(0)=\tanh(a\cdot 0)=0$. Hence, $x=0$ is the unique fixed point.
Let us consider the case $a>1$. For $0<x\ll 1$, $\tanh(ax)-x\approx(a-1)x$. Since $a>1,\tanh(ax)-x>0$. On the other hand, since $|\tanh(ax)|<1$ for all $x$,
$\lim_{x\to\infty}[-1-x]\leq\lim_{x\to\infty}[\tanh(ax)-x]\leq\lim_{x\to\infty}[1-x].$
By the squeeze theorem, $\lim_{x\to\infty}[\tanh(ax)-x]=-\infty$. By the intermediate value theorem, therefore, there exists at least one $x>0$ such that $\tanh(ax)=x$. To establish the uniqueness of the positive fixed point, we investigate the derivative $g^{\prime}(x)=a\operatorname{sech}^{2}(ax)-1$. We find the critical points to be $x=\pm\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}})$. It is straightforward to see that $g^{\prime}(x)>0$ in $\left(-\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}}),\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}})\right)$ and $g^{\prime}(x)<0$ in \$\mathbb{R}\backslash\left(-\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}}),\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}})\right)$. i.e. $g(x)=0$ has exactly two fixed points. Because $g(x)$ is an odd function, if $x^{\ast}$ is a solution, then $-x^{\ast}$ is also a solution. Thus, for $a>1$, there exists a unique positive fixed point if $x>0$ and a unique negative fixed point if $x<0$. ∎
A.2 Proof of Lemma 2
Proof.
(1) Since $(\tanh(ax))^{\prime}=a\operatorname{sech}^{2}(ax)<1$ for all $x>0$, it holds that $x_{n+1}=\phi_{a}(x_{n})<x_{n}$ for all $n\in\mathbb{N}$. Thus the sequence $\{x_{n}\}_{n=1}^{\infty}$ is decreasing. Since $x_{n}>0$ for all $n\in\mathbb{N}$, by the monotone convergence theorem, it converges to the fixed point $x^{*}=0$.
(2) Let $x_{0}<\xi_{a}$. Since $\phi^{\prime}(x)$ decreasing for $x\geq 0$, with $\phi^{\prime}(0)>1$ and $\xi_{a}$ is the unique fixed point for $x>0$, it holds that $x_{n}<x_{n+1}<\xi_{a}$ for all $n\in\mathbb{N}$. Thus, by the monotone convergence theorem, the sequence converges to the fixed point $\xi_{a}$. The proof is similar when $x_{0}>\xi_{a}$. ∎
Appendix B Additional experimental results
B.1 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
(a) Same dimension
(b) Varying dimensions
Figure 4: Here, ’STD’ refers to $\sigma_{z}$. (a) shows the PINN loss for the Burgers’ equation, using an FFNN with 30 layers and 32 nodes in each hidden layer. (b) shows the PINN loss for an FFNN with 30 layers, where the hidden layers alternate between 64 and 32 nodes, repeated 15 times. Each experiment was repeated 10 times with different random seeds.
B.2 Width independence in Classification tasks
Figure 5: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the MNIST dataset, with 10 different random seeds.
Figure 6: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the Fashion MNIST dataset, with 10 different random seeds.
Figure 7: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the CIFAR-10 dataset, with 10 different random seeds.